f ( x ) is a linear function, f ( − 5 ) = − 2 , and f ( 5 ) = − 1 , find an equation for f ( x )
Answer #1

Given 2 f (-5) = f(3) = -1 linear function f(x) = mx te 5-(-5) -5m + C = -2 (5) smtc -) -(1) By Eq (1) و ( 20 = - 3 -34 o in